OpenStudy (anonymous):

Which equation is not equivalent to the others? zero point four x equals two point four three x equals one point eight zero point five x equals three x equals six

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

0.4x = 2.4 x = 2.4/0.4 x = 6 3x = 1.8 x = 1.8/3 x = 0.6 0.5x = 3 x = 3/0.5 x = 6 x = 6 odd man out is 3x = 1.8 :)